OpenStudy (anonymous):

From 1980 through 1990, the population of Country X increased by 100 percent. From 1990 to 2000, The population increased by 50 percent. What was the combined increase for the period 1980-2000 A. 150% B. 166 2/3% C.175% D.200% Would some one please break this down for me cause I thought it was A but the book has D.

OpenStudy (amistre64):

50% is from the value in 1990, not the value in 1980

OpenStudy (amistre64):

lets try it with numbers to see: 5 increased by 100% = 5+5 = 10 10 increased by 50% = 10+5 = 15 5 +100% + 100% = 15

OpenStudy (anonymous):

so then it would be 2300 percent oh ok now i see